### v4

# The Pell sequence and cyclotomic matrices involving squares over finite fields

Abstract.
In this paper, by some arithmetic properties of the Pell sequence and some $p$-adic tools, we study certain cyclotomic matrices involving squares over finite fields. For example, let $1=s_{1},s_{2},\cdots,s_{(q-1)/2}$ be all the nonzero squares over $\mathbb{F}_{q}$, where $q=p^{f}$ is an odd prime power with $q\geq 7$. We prove that the matrix
$B_{q}((q-3)/2)=\left[\left(s_{i}+s_{j}\right)^{(q-3)/2}\right]_{2\leq i,j\leq(q-1)/2}$
is a singular matrix whenever $f\geq 2$. Also, for the case $q=p$, we show that
$\det B_{p}((p-3)/2)=0\Leftrightarrow Q_{p}\equiv 2\ ({\rm{mod}}\ p^{2}\mathbb{Z}),$
where $Q_{p}$ is the $p$-th term of the companion Pell sequence $\{Q_{i}\}_{i=0}^{\infty}$ defined by $Q_{0}=Q_{1}=2$ and $Q_{i+1}=2Q_{i}+Q_{i-1}$.
Key words and phrases:
the Pell sequence, cyclotomic matrices, character sums over finite fields.
2020 Mathematics Subject Classification. Primary 11L05, 15A15; Secondary 11R18, 12E20.
This research was supported by the Natural Science Foundation of China (Grant Nos. 12101321, 12201291 and 12371004).
*Corresponding author
1. Introduction
$1.$ Notation
In this paper, the symbol $p$ always denotes an odd prime. We let $\mathbb{Q}_{p}$ be the $p$-adic number field and let $\mathbb{Z}_{p}$ be the ring of all $p$-adic integers. As usual, we use $\mathbb{C}_{p}$ to denote the completion of the algebraic closure of $\mathbb{Q}_{p}$. In addition, we let
${\rm ord}_{p}:\ \mathbb{C}_{p}\rightarrow\mathbb{R}$
be the $p$-adic order function over $\mathbb{C}_{p}$.
Let $q=p^{f}$ be an odd prime power with $f\in\mathbb{Z}^{+}$. The symbol $\mathbb{F}_{q}$ denotes the finite field with $q$ elements and $\mathbb{F}_{q}^{\times}=\mathbb{F}_{q}\setminus\{0\}$ is the multiplicative cyclic group of all nonzero elements over $\mathbb{F}_{q}$. The group of all multiplicative characters of $\mathbb{F}_{q}$ is denoted by $\widehat{\mathbb{F}_{q}^{\times}}$, and the trivial character is written by $\varepsilon$. Throughout this paper, for any multiplicative character $\psi:\mathbb{F}_{q}^{\times}\rightarrow\mathbb{C}$ (or $\mathbb{C}_{p}$), we extend $\psi$ to $\mathbb{F}_{q}$ by letting $\psi(0)=0$.
Let $\zeta_{p}$ be a primitive $p$-th root of unity over $\mathbb{C}$ (or $\mathbb{C}_{p}$). Then for any $\psi\in\widehat{\mathbb{F}_{q}^{\times}}$, the Gauss sum of $\psi$ is defined by
$G_{q}(\psi)=\sum_{x\in\mathbb{F}_{q}}\psi(x)\zeta_{p}^{{\rm Tr}_{\mathbb{F}_{q}/\mathbb{F}_{p}}(x)},$
where ${\rm Tr}_{\mathbb{F}_{q}/\mathbb{F}_{p}}:\mathbb{F}_{q}\rightarrow\mathbb{F}_{p}$ is the trace map. Also, for any $\psi,\chi\in\widehat{\mathbb{F}_{q}^{\times}}$, the Jacobi sum of $\psi$ and $\chi$ is defined by
$J_{q}(\psi,\chi)=\sum_{x\in\mathbb{F}_{q}}\psi(x)\chi(1-x).$
For any positive integer $a$ and nonnegative integer $b$, the symbol $\binom{a}{b}$ is the binomial coefficient, that is,
$\binom{a}{b}=\frac{a(a-1)\cdots(a-b+1)}{b!}.$
Finally, for any square matrix $M$ over a field, $\det M$ denotes the determinant of $M$.
1.2. Background and motivation
The investigations of cyclotomic matrices involving elements over finite fields were initiated by Lehmer [8] and Carlitz [1]. For instance, for any nontrivial multiplicative character $\psi$ modulo $p$, Carlitz [1] first studied the cyclotomic matrices
(1.1) $C_{p}^{-}(\psi):=\left[\psi(j-i)\right]_{1\leq i,j\leq p-1},$
and
(1.2) $C_{p}^{+}(\psi):=\left[\psi(j+i)\right]_{1\leq i,j\leq p-1}.$
Carlitz [1, Theorem 4 and Theorem 5] proved that
$\det C_{p}^{-}(\psi)=(-1)^{(p-1)/m}\cdot\frac{G_{p}(\psi)^{p-1}}{p},$
and
$\det C_{p}^{+}(\psi)=\begin{cases}\frac{1}{p}(-1)^{\frac{p-1}{2m}}G_{p}(\psi)^{p-1}&\mbox{if}\ m\equiv 1\ ({\rm{mod}}\ 2),\\ \frac{1}{p}(-1)^{\frac{p-1}{m}}\delta(\psi)^{p-1}G_{p}(\psi)^{p-1}&\mbox{if}\ m\equiv 0\ ({\rm{mod}}\ 2),\end{cases}$
where $m=\min\{k\in\mathbb{Z}^{+}:\psi^{k}=\varepsilon\}$ is the order of $\psi$ and
$\delta(\psi)=\begin{cases}1&\mbox{if}\ \psi(-1)=1,\\ -{\bf i}&\mbox{if}\ \psi(-1)=-1.\end{cases}$
Along Carlitz’s work, Chapman [2] further studied some variants of the matrix $C_{p}^{+}(\psi)$. In fact, let $(\frac{\cdot}{p})$ denotes the Legendre symbol, i.e., the unique quadratic multiplicative character of $\mathbb{F}_{p}$. Chapman considered the matrices
$C_{p}^{(0)}:=\left[\left(\frac{i+j}{p}\right)\right]_{0\leq i,j\leq(p-1)/2},$
and
$C_{p}^{(1)}:=\left[\left(\frac{i+j}{p}\right)\right]_{1\leq i,j\leq(p-1)/2}.$
Although Chapman only changed the sizes of the matrices, the calculations of $\det C_{p}^{(0)}$ and $\det C_{p}^{(1)}$ became extremely complicated. In fact, for $p\equiv 1\ ({\rm{mod}}\ 4)$, let $\epsilon_{p}>1$ and $h_{p}$ denote the fundamental unit and class number of $\mathbb{Q}(\sqrt{p})$ respectively, and write
$\epsilon_{p}^{h_{p}}=a_{p}+b_{p}\sqrt{p}$
with $a_{p},b_{p}\in\mathbb{Q}$. Then Chapman [2] showed that
$\det C_{p}^{(0)}=\begin{cases}(-1)^{(p+3)/4}2^{(p-1)/2}a_{p}&\mbox{if}\ p\equiv 1\ ({\rm{mod}}\ 4),\\ -2^{(p-1)/2}&\mbox{if}\ p\equiv 3\ ({\rm{mod}}\ 4),\end{cases}$
and that
$\det C_{p}^{(1)}=\begin{cases}(-1)^{(p-1)/4}2^{(p-1)/2}b_{p}&\mbox{if}\ p\equiv 1\ ({\rm{mod}}\ 4),\\ 0&\mbox{if}\ p\equiv 3\ ({\rm{mod}}\ 4).\end{cases}$
Chapman also investigated some variants of $C_{p}^{-}(\psi)$, and the most well-known variant among them is Chapman’s “evil determinant”:
$\det C_{p}^{-}:=\det\left[\left(\frac{j-i}{p}\right)\right]_{1\leq i,j\leq(p+1)/2}.$
Let
$\epsilon_{p}^{(2+(-1)^{(p+3)/4})h_{p}}=a_{p}^{\prime}+b_{p}^{\prime}\sqrt{p}$
with $a_{p}^{\prime},b_{p}^{\prime}\in\mathbb{Q}$. Then Chapman conjectured that
$\det C_{p}^{-}=\begin{cases}-a_{p}^{\prime}&\mbox{if}\ p\equiv 1\ ({\rm{mod}}\ 4),\\ 1&\mbox{if}\ p\equiv 3\ ({\rm{mod}}\ 4).\end{cases}$
Vsemirnov [16, 17] later confirmed this tough conjecture by using ingenious and sophisticated matrix decompositions.
Sun [13] studied the variants of the above matrices from another perspective. For example, Sun [13] consider the matrix
(1.3) $S_{p}:=\left[\left(\frac{i^{2}+j^{2}}{p}\right)\right]_{1\leq i,j\leq(p-1)/2}$
which involves the nonzero squares over $\mathbb{F}_{p}$. Sun [13, Theorem 1.2] proved that $\det S_{p}\ {\rm mod}\ p\mathbb{Z}$ is a nonzero square over $\mathbb{Z}/p\mathbb{Z}=\mathbb{F}_{p}$. Moreover, Sun conjectured that $\det S_{p}$ is indeed a square of some integer whenever $p\equiv 3\ ({\rm{mod}}\ 4)$. This conjecture was later confirmed by Alekseyev and Krachun. For the case $p\equiv 1\ ({\rm{mod}}\ 4)$, if we write $p=a^{2}+4b^{2}$ with $a,b\in\mathbb{Z}$ and $a\equiv 1\ ({\rm{mod}}\ 4)$, then Cohen, Sun and Vsemirnov conjectured that $S_{p}/a$ is also a square of certain integer. This conjecture was later confirmed by the first author [18].
Recently, for any positive integer $m$, Sun [14] considered the matrix
(1.4) $S_{p}(m):=\left[\left(i^{2}+j^{2}\right)^{m}\right]_{1\leq i,j\leq(p-1)/2},$
and obtained several interesting results concerning $\det S_{p}(m)$. For example, when $p\equiv 3\ ({\rm{mod}}\ 4)$, Sun [14, Theorem 1.2(ii)] proved that
$\det S_{p}(p-3)=\det\left[\frac{1}{(i^{2}+j^{2})^{2}}\right]_{1\leq i,j\leq(p-1)/2}\equiv\frac{1}{4}\prod_{r=1}^{(p-3)/4}\left(r+\frac{1}{4}\right)^{2}\ ({\rm{mod}}\ p\mathbb{Z}_{p}).$
After introducing the above relevant research results, we now state our research motivations. Inspired by Sun’s matrix $S_{p}(m)$ defined by Eq. (1.4), in this paper, we mainly consider a variant of $S_{p}(m)$. In fact, let $q=2n+1$ and let
$\left\{s_{1}=1,s_{2},\cdots,s_{n}\right\}=\{x^{2}:x\in\mathbb{F}_{q}^{\times}\}$
be the set of all nonzero squares over $\mathbb{F}_{q}$. Then we define the matrix
(1.5) $B_{q}(m):=\left[\left(s_{i}+s_{j}\right)^{m}\right]_{2\leq i,j\leq n}.$
Before studying the properties of $B_{q}(m)$, let’s first have a brief discussion. Let
$h(t)=a_{m-1}t^{m-1}+\cdots+a_{1}t+a_{0}$
be a polynomial over a commutative ring with $\deg(h(t))\leq m-1$. Then it is known that (see [6, Lemma 9])
(1.6) $\det\left[h(x_{i}+y_{j})\right]_{1\leq i,j\leq m}=a_{m-1}^{m}\cdot\prod_{r=0}^{m-1}\binom{m-1}{r}\cdot\prod_{1\leq i<j\leq m}\left(x_{i}-x_{j}\right)\left(y_{j}-y_{i}\right).$
By this we immediately see that $B_{q}(m)$ is singular whenever $m\leq n-3$. Hence it is meaningful to consider the cases of $n-2\leq m\leq q-1$. In this paper, we focus on the cases of m=n2,n1,nm=n-2,n-1,n, and we will see later that the methods used in each case are very different.
1.3. The Pell sequence
Before the statements of our main results, we introduce the Pell sequence. The Pell sequence is an infinite sequence of integers that comprise the denominators of the closest rational approximations to the square root of 2. Specifically, the Pell sequence $\{P_{i}\}_{i=0}^{\infty}$ is defined by
$P_{0}=0,P_{1}=1,P_{i+1}=2P_{i}+P_{i-1},$
and its companion sequence $\{Q_{i}\}_{i=0}^{\infty}$ is defined by
$Q_{0}=Q_{1}=2,Q_{i+1}=2Q_{i}+Q_{i-1}.$
For any integer $i\geq 0$, it is also known that
$P_{i}=\frac{1}{2\sqrt{2}}\left((1+\sqrt{2})^{i}-(1-\sqrt{2})^{i}\right)=\sum_{k=0}^{\lfloor(i-1)/2\rfloor}\binom{i}{2k+1}2^{k},$
and
$Q_{i}=(1+\sqrt{2})^{i}+(1-\sqrt{2})^{i}.$
By this, it is easy to see that
$P_{p}\equiv 2^{(p-1)/2}\equiv\left(\frac{2}{p}\right)\ ({\rm{mod}}\ p\mathbb{Z}),$
and that
$Q_{p}\equiv 2\ ({\rm{mod}}\ p\mathbb{Z}).$
1.4. Main results
Now we state our first result, which involves the cases m=n1,n2m=n-1,n-2.
Theorem 1.1.
Let $q=p^{f}=2n+1\geq 7$ be an odd prime power with $p$ prime and $f\in\mathbb{Z}^{+}$. Then the following results hold.
(i) $B_{q}(n-1)$ is a singular matrix whenever $f\geq 2$. Moreover, if $f=1$, then
$\det B_{p}(n-1)=\frac{2\cdot((n-1)!)^{n}}{(0!1!\cdots(n-1)!)^{2}}\cdot a_{p}\in\mathbb{F}_{p},$
where
$a_{p}=\frac{2-Q_{p}}{p}\ {\rm mod}\ p\mathbb{Z}\in\mathbb{Z}/p\mathbb{Z}=\mathbb{F}_{p}.$
(ii) $B_{q}(n-2)$ is singular if and only if the integer $f\geq 2$. When $f=1$, we have
$\det B_{p}(n-2)=\left(-\frac{1}{2}\right)^{n-2}\cdot\frac{((n-2)!)^{n-1}}{(0!1!\cdots(n-2)!)^{2}}\in\mathbb{F}_{p}^{\times}.$
By Theorem 1.1, we obtain the following result.
Corollary 1.1.
Let $p=2n+1\geq 7$ be an odd prime. Then the following results hold.
(i) $B_{p}(n-1)$ is singular if and only if
$Q_{p}\equiv 2\ ({\rm{mod}}\ p^{2}\mathbb{Z}).$
Moreover, if $p\equiv 1\ ({\rm{mod}}\ 4)$ , then
$\left(\frac{\det B_{p}(n-1)}{p}\right)=\left(\frac{2a_{p}}{p}\right).$
If $p\equiv 3\ ({\rm{mod}}\ 4)$, then
$\left(\frac{\det B_{p}(n-1)}{p}\right)=(-1)^{\frac{h(-p)-1}{2}}\left(\frac{a_{p}}{p}\right),$
where h(p)h(-p) is the class number of the imaginary quadratic field $\mathbb{Q}(\sqrt{-p})$.
(ii) Suppose $p\equiv 1\ ({\rm{mod}}\ 4)$. Then
$\left(\frac{\det B_{p}(n-2)}{p}\right)=\left(\frac{6}{p}\right).$
Suppose $p\equiv 3\ ({\rm{mod}}\ 4)$. Then
$\left(\frac{\det B_{p}(n-2)}{p}\right)=\left(\frac{-2}{p}\right).$
Remark $1.$
With the help of a computer and by the referee’s comments, we verify that
(1.7) $\left\{p:\ p\ \text{is a prime with}\ 7\leq p\leq 10^{6}\ \text{and}\ Q_{p}\equiv 2\ ({\rm{mod}}\ p^{2}\mathbb{Z})\right\}=\left\{13,31\right\}.$
Motivated by this, we pose the following conjecture.
Conjecture $1.$
Let $p\geq 7$ be an odd prime. Then $B_{p}(n-1)$ is singular if and only if $p\in\{13,31\}$.
We next turn to the case $m=n$.
Theorem 1.2.
Let $q=p^{f}=2n+1\geq 7$ be an odd prime power with $p$ prime and $f\in\mathbb{Z}^{+}$. Then $B_{q}(n)$ is singular whenever $f\geq 2$. Moreover, if $f=1$, then
$\det B_{p}(n)=(-1)^{n}\left(\frac{1}{2}\right)^{n-2}\cdot\frac{(n!)^{n+1}}{(0!1!\cdots n!)^{2}}\cdot b_{p}\in\mathbb{F}_{p},$
where
$b_{p}=\frac{2(\frac{2}{p})-2P_{p}-p}{p}\ {\rm mod}\ p\mathbb{Z}\in\mathbb{Z}/p\mathbb{Z}=\mathbb{F}_{p}.$
Similar to Corollary 1.1, we have the following result.
Corollary 1.2.
Let $p=2n+1\geq 7$ be a prime. Then
$\left(\frac{\det B_{p}(n)}{p}\right)=\left(\frac{-2b_{p}}{p}\right),$
and hence $B_{p}(n)$ is singular if and only if
$2P_{p}\equiv 2\left(\frac{2}{p}\right)-p\ ({\rm{mod}}\ p^{2}\mathbb{Z}).$
Remark 1.2.
As (1.7), it is natural to consider the odd prime $p$, which satisfies the congruence
$2P_{p}\equiv 2\left(\frac{2}{p}\right)-p\ ({\rm{mod}}\ p^{2}\mathbb{Z}).$
By the referee’s comments, we have
(1.8) $\left\{p:\ p\ \text{is a prime with}\ 7\leq p\leq 10^{6}\ \text{and}\ 2P_{p}\equiv 2\left(\frac{2}{p}\right)-p\ ({\rm{mod}}\ p^{2}\mathbb{Z})\right\}=\left\{29\right\}.$
Using essentially the same method appeared in the proof of Theorem 1.2, we can obtain some satisfactory results on certain variants of Carlitz’s matrice $C_{p}^{-}(\psi)$ and $C_{p}^{+}(\psi)$ defined by Eq. (1.1) and Eq. (1.2) respectively.
In fact, let $\mathbb{F}_{q}^{\times}=\{x_{1}=1,x_{2},\cdots,x_{q-1}\}$. For any nontrivial character $\psi\in\widehat{\mathbb{F}_{q}^{\times}}$, we define
$D_{q}^{-}(\psi):=\left[\psi(x_{j}-x_{i})\right]_{2\leq i,j\leq q-1},$
and
$D_{q}^{+}(\psi):=\left[\psi(x_{j}+x_{i})\right]_{2\leq i,j\leq q-1}.$
We have the following result.
Theorem 1.3.
Let $q=p^{f}\geq 3$ be an odd prime power. Then for any nontrivial character $\psi\in\widehat{\mathbb{F}_{q}^{\times}}$, we have
$\det D_{q}^{-}(\psi)=-\frac{1+\psi(-1)}{q^{2}}G_{q}(\psi)^{q-1},$
and
$\det D_{q}^{+}(\psi)=\frac{(-1)^{(q+1)/2}\cdot\psi(-1)}{q^{2}}\left(2-\overline{\psi(2)}\right)G_{q}(\psi)^{q-1}.$
Let $\psi=\phi$ be the unique quadratic character of $\mathbb{F}_{q}$, i.e.,
$\phi(x)=\begin{cases}1&\mbox{if}\ x\in\{s_{1},s_{2},\cdots,s_{(q-1)/2}\},\\ 0&\mbox{if}\ x=0,\\ -1&\mbox{otherwise.}\end{cases}$
Then $D_{p}^{-}(\phi)$ is an integer matrices. Note that (see [3, Corollary 3.7.6])
[…]
By this and Theorem 1.3, we can obtain the following corollary directly.
Corollary 1.3.
Let $q=p^{f}\geq 3$ be an odd prime power. Then $D_{q}^{-}(\phi)$ is singular if and only if $q\equiv 3\ ({\rm{mod}}\ 4)$. For $q\equiv 1\ ({\rm{mod}}\ 4)$, we have
$\det D_{q}^{-}(\phi)=-2q^{(q-5)/2}.$
1.5. Outline of this paper
We will prove Theorem 1.1 and its corollary in Section 2. In Section 3, we shall introduce some necessary results on almost circulant matrices and some $p$-adic tools. The proofs of Theorems 1.2 and 1.3 will be given in Section 4 and Section 5 respectively.
2. Proofs of Theorem 1.1 and its corollary
Recall that n=(q1)/2n=(q-1)/2 and $s_{1}=1,s_{2},\cdots,s_{n}$ are exactly all the nonzero squares over $\mathbb{F}_{q}$. Also, for any integers $a,b$ with $a\leq b$, we use the symbol $[a,b]$ to denote the set $\{a,a+1,\cdots,b\}$.
We begin with the following result.
Lemma 2.1.
Let $q=p^{f}=2n+1\geq 5$ be an odd prime power with $p$ prime and $f\in\mathbb{Z}^{+}$. Then
$(-1)^{\frac{(n-1)(n-2)}{2}}\prod_{2\leq i<j\leq n}\left(s_{j}-s_{i}\right)^{2}=\left(-\frac{1}{2}\right)^{n-2}\in\mathbb{F}_{p}.$
Proof.
Let
$F(t)=\frac{t^{n}-1}{t-1}=t^{n-1}+t^{n-2}+\cdots+t+1$
be a polynomial over $\mathbb{F}_{q}$. Note that $x$ is a nonzero square over $\mathbb{F}_{q}$ if and only if $x^{n}=1$. Hence it is easy to verify that
(2.1) $F(t)=\frac{t^{n}-1}{t-1}=t^{n-1}+t^{n-2}+\cdots+t+1=\prod_{2\leq j\leq n}\left(t-s_{j}\right).$
By Eq. (2.1) we obtain that $(-1)^{n-1}\prod_{2\leq j\leq n}s_{j}$ is equal to the constant term of $F(t)$, i.e.,
(2.2) $\prod_{2\leq j\leq n}s_{j}=(-1)^{n-1}.$
Also, clearly $(-1)^{n-1}\prod_{2\leq j\leq n}(s_{j}-1)$ is equal to the constant term of $F(t+1)$, that is,
(2.3) $\prod_{2\leq j\leq n}(s_{j}-1)=(-1)^{n-1}\cdot n.$
Next we consider the product
$(-1)^{\frac{(n-1)(n-2)}{2}}\prod_{2\leq i<j\leq n}\left(s_{j}-s_{i}\right)^{2}.$
Let $F^{\prime}(t)$ be the formal derivative of $F(t)$. Then one can verify that
$(-1)^{\frac{(n-1)(n-2)}{2}}\prod_{2\leq i<j\leq n}\left(s_{j}-s_{i}\right)^{2}$ $=\prod_{2\leq i\neq j\leq n}\left(s_{j}-s_{i}\right)$
$=\prod_{2\leq j\leq n}\prod_{i\in[2,n]\setminus\{j\}}\left(s_{j}-s_{i}\right)$
(2.4) $=\prod_{2\leq j\leq n}F^{\prime}(s_{j}).$
As $(t-1)F(t)=t^{n}-1$, we have
$F(t)+(t-1)F^{\prime}(t)=nt^{n-1},$
and hence
$F^{\prime}(s_{j})=\frac{ns_{j}^{n-1}}{s_{j}-1}=\frac{n}{s_{j}(s_{j}-1)}$
for any $j\in[2,n]$. By the above results,
(2.5) $(-1)^{\frac{(n-1)(n-2)}{2}}\prod_{2\leq i<j\leq n}\left(s_{j}-s_{i}\right)^{2}=\prod_{2\leq j\leq n}\frac{n}{s_{j}(s_{j}-1)}.$
Combining Eq. (2.5) with Eq. (2.2) and Eq. (2.3), we obtain
$(-1)^{\frac{(n-1)(n-2)}{2}}\prod_{2\leq i<j\leq n}\left(s_{j}-s_{i}\right)^{2}=n^{n-2}=\left(-\frac{1}{2}\right)^{n-2}\in\mathbb{F}_{p}.$
This completes the proof. ∎
Before the statement of next lemma, we introduce the following notations. Let $l$ be a positive integer and let $t_{1},t_{2},\cdots,t_{l}$ be variables. Then for any $k\in[1,l]$, the $k$-th elementary symmetric polynomial of $t_{1},t_{2},\cdots,t_{l}$ is defined by
$\sigma_{k}(t_{1},t_{2},\cdots,t_{l})=\sum_{1\leq i_{1}<i_{2}<\cdots<i_{k}\leq l}\prod_{1\leq j\leq k}t_{i_{j}}.$
In addition, we let
$\sigma_{0}(t_{1},t_{2},\cdots,t_{l})=1.$
In 2022, Grinberg, Sun and Zhao [4, Theorem 3.1] obtained the following result.
Lemma 2.2.
Let $l$ be a positive integer. Then
(2.6) $\det\left[\left(x_{i}+y_{j}\right)^{l}\right]_{1\leq i,j\leq l}=(-1)^{\frac{l(l-1)}{2}}\cdot\prod_{1\leq i<j\leq l}\left(x_{j}-x_{i}\right)\left(y_{j}-y_{i}\right)\cdot\sum_{k=0}^{l}U_{k},$
where
$U_{k}=\sigma_{k}\left(x_{1},\cdots,x_{l}\right)\sigma_{l-k}\left(y_{1},\cdots,y_{l}\right)\cdot\prod_{r\in[0,l]\setminus\{k\}}\binom{l}{r}$.
We now need the following results related to some congruences involving the Pell sequence, which were obtained by Z.-H. Sun [10, Theorem 4.1], Z.-W. Sun [11, Final Remark] and Z.-W. Sun [12, Remark 3.1] respectively.
Lemma 2.3.
Let $p\geq 7$ be a prime. Then the following congruences hold.
(2.7) $\sum_{k=1}^{p-1}\frac{2^{k}}{k}\equiv\frac{2-2^{p}}{p}-\frac{7}{12}p^{2}B_{p-3}\ ({\rm{mod}}\ p^{3}\mathbb{Z}_{p}),$
where $B_{p-3}$ is the (p3)(p-3)-th Bernoulli number.
(2.8) $\sum_{k=1}^{\frac{p-1}{2}}\frac{1}{k\cdot 2^{k}}\equiv-2^{\frac{p+1}{2}}\cdot\frac{P_{p}-2^{\frac{p-1}{2}}}{p}\ ({\rm{mod}}\ p\mathbb{Z}_{p}).$
(2.9) $4\left(\frac{2}{p}\right)P_{p}\equiv 2+Q_{p}\ ({\rm{mod}}\ p^{2}\mathbb{Z}).$
Now we are in a position to prove Theorem 1.1.
Proof of Theorem 1.1. (i) By Lemma 2.2 and Lemma 2.1 we obtain
(2.10) $\det B_{q}(n-1)=(-1)^{\frac{(n-1)(n-2)}{2}}\prod_{2\leq i<j\leq n}\left(s_{j}-s_{i}\right)^{2}\cdot\sum_{k=0}^{n-1}W_{k}=\left(-\frac{1}{2}\right)^{n-2}\cdot\sum_{k=0}^{n-1}W_{k},$
where
(2.11) $W_{k}=\sigma_{k}\left(s_{2},\cdots,s_{n}\right)\sigma_{n-1-k}\left(s_{2},\cdots,s_{n}\right)\cdot\prod_{r\in[0,n-1]\setminus\{k\}}\binom{n-1}{r}$
for any $k\in[0,n-1]$.
We next consider $W_{k}$. By Eq. (2.1) we see that
$F(t)=t^{n-1}+t^{n-2}+\cdots+t+1=\prod_{2\leq j\leq n}\left(t-s_{j}\right)=\sum_{k=0}^{n-1}(-1)^{k}\sigma_{k}\left(s_{2},\cdots,s_{n}\right)t^{n-1-k},$
and hence for any $k\in[0,n-1]$ we have
(2.12) $\sigma_{k}\left(s_{2},\cdots,s_{n}\right)=(-1)^{k}.$
By Eq. (2.11) and Eq. (2.12) we obtain
(2.13) $W_{k}=(-1)^{n-1}\prod_{r\in[0,n-1]\setminus\{k\}}\binom{n-1}{r}$.
Suppose that $f\geq 2$. As $q=p^{f}\geq 9$, we have
$n-1=\frac{q-3}{2}=\frac{p-3}{2}\cdot 1+\frac{p-1}{2}\cdot p+\cdots+\frac{p-1}{2}\cdot p^{f-1}>\frac{p+1}{2}.$
By the Lucas congruence, for $j\in\{(p-1)/2,(p+1)/2\}$, we have
$\binom{n-1}{j}\equiv\binom{(p-3)/2}{j}\binom{(p-1)/2}{0}\cdots\binom{(p-1)/2}{0}\equiv 0\ ({\rm{mod}}\ p\mathbb{Z}).$
Thus,
$\left|\left\{r\in[0,n-1]:\ \binom{n-1}{r}\equiv 0\ ({\rm{mod}}\ p\mathbb{Z})\right\}\right|\geq 2.$
This, together with Eq. (2.13), implies that $W_{k}=0$ (as an element of $\mathbb{F}_{q}$) if $f\geq 2$. Hence by Eq. (2.10) we see that $B_{q}(n-1)$ is singular whenever $f\geq 2$.
Suppose now $f=1$. Then for any $r\in[0,n-1]$ we have
$\binom{n-1}{r}\not\equiv 0\ ({\rm{mod}}\ p\mathbb{Z}).$
Hence by Eq. (2.13) we obtain
$\sum_{k=0}^{n-1}W_{k}$ $=(-1)^{n-1}\left(\sum_{r=0}^{n-1}\binom{n-1}{r}^{-1}\right)\cdot\prod_{r\in[0,n-1]}\binom{n-1}{r}$
(2.14) $=(-1)^{n-1}\cdot\frac{n}{2^{n}}\left(\sum_{k=1}^{n}\frac{2^{k}}{k}\right)\cdot\frac{((n-1)!)^{n}}{(0!1!\cdots(n-1)!)^{2}},$
where the last equality follows from Sury’s identity [15]
$\sum_{r=0}^{n-1}\binom{n-1}{r}^{-1}=\frac{n}{2^{n}}\sum_{k=1}^{n}\frac{2^{k}}{k}$
and
$\prod_{r\in[0,n-1]}\binom{n-1}{r}=\frac{((n-1)!)^{n}}{(0!1!\cdots(n-1)!)^{2}}.$
We next consider
$\sum_{k=1}^{n}\frac{2^{k}}{k}\ {\rm mod}\ p\mathbb{Z}_{p}.$
One can verify that
$\sum_{k=1}^{p-1}\frac{2^{k}}{k}$ $=\sum_{k=1}^{n}\frac{2^{k}}{k}+\sum_{k=1}^{n}\frac{2^{p-k}}{p-k}$
$\equiv\sum_{k=1}^{n}\frac{2^{k}}{k}-2\sum_{k=1}^{n}\frac{1}{k\cdot 2^{k}}\ ({\rm{mod}}\ p\mathbb{Z}_{p}).$
Combining this with Eq. (2.7) and Eq. (2.8), we obtain
$\sum_{k=1}^{n}\frac{2^{k}}{k}$ $\equiv\frac{2-2^{p}}{p}-2^{\frac{p+3}{2}}\frac{P_{p}-2^{\frac{p-1}{2}}}{p}$
$\equiv\frac{2-2^{p}}{p}-4\left(\frac{2}{p}\right)\frac{P_{p}-2^{\frac{p-1}{2}}}{p}\ ({\rm{mod}}\ p\mathbb{Z}_{p}).$
By this and Eq. (2.9), we obtain
(2.15) $\sum_{k=1}^{n}\frac{2^{k}}{k}\equiv\frac{-Q_{p}-2^{p}+(\frac{2}{p})2^{\frac{p+3}{2}}}{p}\equiv\frac{2-Q_{p}}{p}\ ({\rm{mod}}\ p\mathbb{Z}_{p}),$
where the last congruence follows from
$2-\left(-2^{p}+\left(\frac{2}{p}\right)2^{\frac{p+3}{2}}\right)=2\left(2^{\frac{p-1}{2}}-\left(\frac{2}{p}\right)\right)^{2}\equiv 0\ ({\rm{mod}}\ p^{2}\mathbb{Z}).$
Now combining Eq. (2.15) with Eq. (2) and Eq. (2.10), one can verify that
$\det B_{p}(n-1)=\frac{2\cdot((n-1)!)^{n}}{(0!1!\cdots(n-1)!)^{2}}\cdot a_{p}\in\mathbb{F}_{p},$
where
$a_{p}=\frac{2-Q_{p}}{p}\ {\rm mod}\ p\mathbb{Z}\in\mathbb{Z}/p\mathbb{Z}=\mathbb{F}_{p}.$
This completes the proof of (i).
(ii) Let $h(t)=t^{n-2}$. Then by Eq. (1.6) and Lemma 2.1 we obtain
$\det B_{q}(n-2)$ $=\prod_{r=0}^{n-2}\binom{n-2}{r}\cdot\prod_{2\leq i<j\leq n}\left(-\left(s_{j}-s_{i}\right)^{2}\right)$
$=\prod_{r=0}^{n-2}\binom{n-2}{r}\cdot(-1)^{\frac{(n-1)(n-2)}{2}}\prod_{2\leq i<j\leq n}\left(s_{j}-s_{i}\right)^{2}$
(2.16) $=\left(-\frac{1}{2}\right)^{n-2}\cdot\prod_{r=0}^{n-2}\binom{n-2}{r}$.
Suppose first that $f\geq 2$. Noting that $q\geq 9$, we obtain
$n-2=\frac{q-5}{2}=\frac{p-5}{2}\cdot 1+\frac{p-1}{2}\cdot p+\cdots+\frac{p-1}{2}\cdot p^{f-1}>\frac{p-1}{2}.$
By the Lucas congruence again, we clearly have
$\binom{n-2}{(p-1)/2}\equiv\binom{(p-5)/2}{(p-1)/2}\binom{(p-1)/2}{0}\cdots\binom{(p-1)/2}{0}\equiv 0\ ({\rm{mod}}\ p\mathbb{Z}).$
By this and Eq. (2), we see that $\det B_{q}(n-2)=0$ whenever $f\geq 2$.
Suppose now $f=1$. Then $\binom{n-2}{r}\not\equiv 0\ ({\rm{mod}}\ p\mathbb{Z})$ for any $r\in[0,n-2]$. Hence by Eq. (2) again we obtain
$\det B_{p}(n-2)=\left(-\frac{1}{2}\right)^{n-2}\cdot\frac{((n-2)!)^{n-1}}{(0!1!\cdots(n-2)!)^{2}}\in\mathbb{F}_{p}^{\times}.$
In view of the above, we have completed the proof.∎
We next prove Corollary 1.1.
Proof of Corollary 1.1. (i) If $p\equiv 1\ ({\rm{mod}}\ 4)$, then $n$ is even. Thus,
$\left(\frac{\det B_{p}(n-1)}{p}\right)=\left(\frac{2a_{p}}{p}\right)\left(\frac{(n-1)!}{p}\right)^{n}=\left(\frac{2a_{p}}{p}\right).$
Suppose now $p\equiv 3\ ({\rm{mod}}\ 4)$ and $p>3$. Then $n$ is odd and hence
$\left(\frac{\det B_{p}(n-1)}{p}\right)=\left(\frac{2a_{p}}{p}\right)\left(\frac{(n-1)!}{p}\right)=\left(\frac{-a_{p}}{p}\right)\left(\frac{n!}{p}\right)=(-1)^{\frac{h(-p)-1}{2}}\left(\frac{a_{p}}{p}\right).$
The last equality follows from the Mordell congruence [9]
$n!=\frac{p-1}{2}!\equiv(-1)^{\frac{h(-p)+1}{2}}\ ({\rm{mod}}\ p\mathbb{Z}),$
where h(p)h(-p) is the class number of the imaginary quadratic field $\mathbb{Q}(\sqrt{-p})$. This completes the proof of (i).
(ii) Suppose $p\equiv 1\ ({\rm{mod}}\ 4)$. Then by the Wilson congruence, one can very that
$1\equiv(p-1)!\equiv(-1)^{n}\cdot(n!)^{2}\equiv(n!)^{2}\ ({\rm{mod}}\ p\mathbb{Z}).$
This implies $n!=\sqrt{-1}$ over $\mathbb{F}_{p}$. Thus,
$\left(\frac{n!}{p}\right)\equiv(n!)^{n}\equiv(-1)^{n/2}\equiv\left(\frac{2}{p}\right)\ ({\rm{mod}}\ p\mathbb{Z}),$
that is,
$\left(\frac{n!}{p}\right)=\left(\frac{2}{p}\right).$
Now by this and Theorem 1.1(ii), and noting that $n$ is even in this case, we obtain
$\left(\frac{\det B_{p}(n-2)}{p}\right)=\left(\frac{(n-2)!}{p}\right)=\left(\frac{n!}{p}\right)\left(\frac{(n-1)n}{p}\right)=\left(\frac{6}{p}\right),$
where the last equality follows from n(n1)=3/4n(n-1)=3/4 over $\mathbb{F}_{p}$.
Suppose now $p\equiv 3\ ({\rm{mod}}\ 4)$. Then $n$ is odd. By Theorem 1.1(ii) we clearly have
$\left(\frac{\det B_{p}(n-2)}{p}\right)=\left(\frac{-2}{p}\right).$
In view of the above, we have completed the proof. ∎
3. Preparations for the proof of Theorem 1.2
3.1. A lemma on almost circulant matrices
Let $n\geq 2$ be an integer and let ${\bm{v}}=(a_{0},a_{1},\cdots,a_{n-1})\in\mathbb{C}^{n}$. The circulant matrix of $\bm{v}$} is an $n\times n$ matrix defined by
$C_{n}({\bm{v}})=[a_{j-i}]_{0\leq i,j\leq n-1},$
where $a_{s}=a_{t}$ whenever $s\equiv t\ ({\rm{mod}}\ n)$, that is,
$C_{n}({\bm{v}})=\left[\begin{array}[]{ccccc}a_{0}&a_{1}&\cdots&a_{n-2}&a_{n-1}\\ a_{n-1}&a_{0}&\cdots&a_{n-3}&a_{n-2}\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ a_{2}&a_{3}&\cdots&a_{0}&a_{1}\\ a_{1}&a_{2}&\cdots&a_{n-1}&a_{0}\end{array}\right].$
Recently, the first author and the second author [19, Theorem 4.1] defined the almost circulant matrix $W_{n}({\bm{v}})$ of $\bm{v}$} by
$W_{n}({\bm{v}})=\left[a_{j-i}\right]_{1\leq i,j\leq n-1},$
and obtained the following result.
Lemma 3.1.
Let $\lambda_{0},\lambda_{1},\cdots,\lambda_{n-1}$ be all the eigenvalues of $C_{n}({\bm{v}})$. Then
$\det W_{n}({\bm{v}})=\frac{1}{n}\sum_{l=0}^{n-1}\prod_{k\in[0,n-1]\setminus\{l\}}\lambda_{k}.$
3.2. Some $p$-adic preparations
Let $\pi\in\mathbb{C}_{p}$ with $\pi^{p-1}+p=0$ and let $\zeta_{\pi}\in\mathbb{C}_{p}$ be a primitive $p$-th root of unity with $\zeta_{\pi}\equiv 1+\pi\ ({\rm{mod}}\ \pi^{2})$.
Recall that $q=p^{f}$. Let $\zeta_{q-1}\in\mathbb{C}_{p}$ be a primitive (q1)(q-1)-th root of unity. Then it is known that $\mathbb{Q}_{p}(\zeta_{q-1})/\mathbb{Q}_{p}$ is an unramified extension and $\mathbb{Q}_{p}(\zeta_{q-1}):\mathbb{Q}_{p}]=f$. Hence
$\mathbb{Z}_{p}[\zeta_{q-1}]/\mathfrak{p}\cong\mathbb{F}_{q},$
where $\mathfrak{p}=p\mathbb{Z}_{p}[\zeta_{q-1}]$ is a prime ideal of $\mathbb{Z}[\zeta_{q-1}]$. From now on, we identify $\mathbb{F}_{q}$ with $\mathbb{Z}_{p}[\zeta_{q-1}]/\mathfrak{p}$. The Teichümuller character $\omega_{q}:\mathbb{F}_{q}\rightarrow\mathbb{C}_{p}$ is a multiplicative character of $\mathbb{F}_{q}$ defined by
(3.1) $\omega_{q}(x\ {\rm mod}\ \mathfrak{p})\equiv x\ ({\rm{mod}}\ \mathfrak{p})$
for any $x\in\mathbb{Z}_{p}[\zeta_{q-1}]$. Also, it is easy to verify that $\omega_{q}$ is a generator of $\widehat{\mathbb{F}_{q}^{\times}}$. For any integer $r\in[0,q-2]$, letting
$r=r_{0}\cdot 1+r_{1}\cdot p+\cdots+r_{f-1}\cdot p^{f-1}$
be the decomposition of $r$ in the base $p$, we define
(3.2) $s(r)=\sum_{i=0}^{f-1}r_{i}.$
For $s(r)$, we have the following known result (see [3, Lemma 3.6.7]).
Lemma 3.2.
For any $r\in[0,q-2]$ we have
$s(r)=(p-1)\sum_{i=0}^{f-1}\left\{\frac{rp^{i}}{q-1}\right\},$
where $\{x\}$ denotes the fractional part of a real number $x$.
Recall that $n=(q-1)/2=(p^{f}-1)/2$. Using Lemma 3.2, we obtain the following result.
Lemma $3.$
Suppose $r\in[0,n-1]$. Then
$s(n)+s(n+r)>s(r).$
Proof.
For any $i\in[0,f-1]$, we set
$x_{i}(r):=\left\{\frac{rp^{i}}{q-1}\right\}.$
Then it is easy to verify that
(3.3) $\frac{1}{2}+\left\{\frac{1}{2}+x_{i}(r)\right\}=\begin{cases}x_{i}(r)+1&\mbox{if}\ 0\leq x_{i}(r)<1/2,\\ x_{i}(r)&\mbox{if}\ 1/2\leq x_{i}(r)<1.\end{cases}$
As $0\leq r\leq n-1$, we have $x_{0}(r)<1/2$. Hence by Eq. (3.3) we obtain
$\frac{1}{2}+\left\{\frac{1}{2}+x_{0}(r)\right\}>x_{0}(r).$
By this, Lemma 3.2 and Eq. (3.3), one can verify that
$s(n)+s(n+r)$ $=(p-1)\sum_{i=0}^{f-1}\left(\frac{1}{2}+\left\{\frac{(n+r)p^{i}}{q-1}\right\}\right)$
$=(p-1)\sum_{i=0}^{f-1}\left(\frac{1}{2}+\left\{\frac{p^{i}}{2}+\frac{rp^{i}}{q-1}\right\}\right)$
$=(p-1)\sum_{i=0}^{f-1}\left(\frac{1}{2}+\left\{\frac{p^{i}-1}{2}+\frac{1}{2}+\frac{rp^{i}}{q-1}\right\}\right)$
$=(p-1)\sum_{i=0}^{f-1}\left(\frac{1}{2}+\left\{\frac{1}{2}+x_{i}(r)\right\}\right)$
$>(p-1)\sum_{i=0}^{f-1}x_{i}(r)$
$=s(r).$
This completes the proof. ∎
Recall that $p$ is an odd prime. We next introduce the $p$-adic Gamma function $\Gamma_{p}:\mathbb{Z}_{p}\rightarrow\mathbb{Z}_{p}^{\times}$, where $\mathbb{Z}_{p}^{\times}$ denotes the group of all $p$-adic units. For any $n\in\mathbb{Z}^{+}$, we define
$\Gamma_{p}(n)=(-1)^{n}\prod_{k\in[1,n-1]\cap\mathbb{Z}_{p}^{\times}}k.$
Since $\mathbb{Z}^{+}$ is dense in $\mathbb{Z}_{p}$ and $\mathbb{Z}_{p}^{\times}$ is a closed multiplicative group, the $p$-adic Gamma function $\Gamma_{p}:\mathbb{Z}_{p}\rightarrow\mathbb{Z}_{p}^{\times}$ is defined by
$\Gamma_{p}(x)=\lim_{i\rightarrow\infty}\Gamma_{p}(x_{i}),$
where $\{x_{i}\}_{i=1}^{\infty}$ is a sequence of positive integers $x_{i}$ with $\lim_{i\rightarrow\infty}x_{i}=x$.
In this paper, we need following result on $\Gamma_{p}$.
Lemma 3.4.
Suppose that $p\geq 5$ is a prime. Let $n\in\mathbb{Z}^{+}$. Then for any $x,y\in\mathbb{Z}_{p}$ we have
$x\equiv y\ ({\rm{mod}}\ p^{n}\mathbb{Z}_{p})\Rightarrow\Gamma_{p}(x)\equiv\Gamma_{p}(y)\ ({\rm{mod}}\ p^{n}\mathbb{Z}_{p}).$
We conclude this section with the following result, which is known as the Gross-Koblitz formula [5, Theorem 1.7].
Lemma 3.5.
Let notations be as above. For any $r\in[0,q-2]$, let the Gauss sum
$G_{q}(\omega_{q}^{-r})=\sum_{x\in\mathbb{F}_{q}}\omega_{q}^{-r}(x)\zeta_{\pi}^{{\rm Tr}_{\mathbb{F}_{q}/\mathbb{F}_{p}}(x)}.$
Then
$G_{q}(\omega_{q}^{-r})=-\pi^{s(r)}\cdot\prod_{i=0}^{f-1}\Gamma_{p}\left(\left\{\frac{rp^{i}}{q-1}\right\}\right),$
where $\Gamma_{p}$ is the $p$-adic Gamma function and $\{x\}$ is the fractional part of a real number $x$.
4. Proof of Theorem 1.2
Let notations be as in Section 3. In this section, we fix a generator $g$ of the cyclic group $\mathbb{F}_{q}^{\times}$. Recall that n=(q1)/2n=(q-1)/2. For any $i\in[0,n-1]$, we let
$a_{i}=\omega_{q}^{-n}\left(1+g^{2i}\right),$
and let
${\bm{v}}=\left(a_{0},a_{1},\cdots,a_{n-1}\right).$
We begin with the following result.
Lemma 4.1.
For any $r\in[0,n-1]$, let
(4.1) $\lambda_{r}:=\frac{(-1)^{r}}{2}J_{q}\left(\omega_{q}^{-n},\omega_{q}^{-r}\right)+\frac{(-1)^{n+r}}{2}J_{q}\left(\omega_{q}^{-n},\omega_{q}^{-(n+r)}\right).$
Then $\lambda_{0}=-1,\lambda_{1},\cdots,\lambda_{n-1}$ are exactly all the eigenvalues of the circulant matrix $C_{n}({\bm{v}})$.
Proof.
For any $r\in[0,n-1]$, one can verify that
$\sum_{0\leq j\leq n-1}\omega_{q}^{-n}\left(1+g^{2j-2i}\right)\omega_{q}^{-r}\left(g^{2j}\right)$
$=$ $\sum_{0\leq j\leq n-1}\omega_{q}^{-n}\left(1+g^{2j-2i}\right)\omega_{q}^{-r}\left(g^{2j-2i}\right)\omega_{q}^{-r}\left(g^{2i}\right)$
$=$ $\sum_{0\leq j\leq n-1}\omega_{q}^{-n}\left(1+g^{2j}\right)\omega_{q}^{-r}\left(g^{2j}\right)\omega_{q}^{-r}\left(g^{2i}\right).$
This implies that for any $r\in[0,n-1]$,
$C_{n}({\bm{v}}){\bm{u}}_{r}=y_{r}{\bm{u}}_{r},$
where the column vector
${\bm{u}}_{r}=\left(\omega_{q}^{-r}(g^{0}),\omega_{q}^{-r}(g^{2}),\cdots,\omega_{q}^{-r}(g^{2(n-1)})\right)^{T},$
and
$y_{r}=\sum_{0\leq j\leq n-1}\omega_{q}^{-n}\left(1+g^{2j}\right)\omega_{q}^{-r}\left(g^{2j}\right).$
As ${\bm{u}}_{0},{\bm{u}}_{1},\cdots,{\bm{u}}_{n-1}$ are clearly linearly independent over $\mathbb{C}$, the numbers $y_{0},y_{1},\cdots,y_{n-1}$ are all the eigenvalues of $C_{n}({\bm{v}})$. Note that
$\frac{1}{2}\left(\varepsilon(x)+\omega_{q}^{-n}(x)\right)=\begin{cases}1&\mbox{if}\ x\in\{s_{1},s_{2},\cdots,s_{n}\},\\ 0&\mbox{otherwise}.\end{cases}$
Thus, for each $r\in[0,n-1]$, we have
$y_{r}$ $=\sum_{0\leq j\leq n-1}\omega_{q}^{-n}\left(1+g^{2j}\right)\omega_{q}^{-r}\left(g^{2j}\right)$
$=\sum_{1\leq j\leq n}\omega_{q}^{-n}\left(1+s_{j}\right)\omega_{q}^{-r}\left(s_{j}\right)$
$=\frac{1}{2}\sum_{x\in\mathbb{F}_{q}}\left(\varepsilon(x)+\omega_{q}^{-n}(x)\right)\omega_{q}^{-n}\left(1+x\right)\omega_{q}^{-r}\left(x\right)$
$=\frac{(-1)^{r}}{2}\sum_{x\in\mathbb{F}_{q}}\omega_{q}^{-n}\left(1+x\right)\omega_{q}^{-r}\left(-x\right)+\frac{(-1)^{n+r}}{2}\sum_{x\in\mathbb{F}_{q}}\omega_{q}^{-n}\left(1+x\right)\omega_{q}^{-(n+r)}\left(-x\right)$
$=\lambda_{r}.$
For $\lambda_{0}$, one can verify that
$\lambda_{0}$ $=\frac{1}{2}J_{q}\left(\omega_{q}^{-n},\varepsilon\right)+\frac{(-1)^{n}}{2}J_{q}\left(\omega_{q}^{-n},\omega_{q}^{{-n}}\right)$
$=\frac{1}{2}J_{q}\left(\omega_{q}^{-n},\varepsilon\right)+\frac{(-1)^{n}}{2}J_{q}\left(\omega_{q}^{-n},\omega_{q}^{{n}}\right)$
$=-\frac{1}{2}\omega_{q}(1)+\frac{(-1)^{n}}{2}(-\omega_{q}^{-n}(-1))$
$=-1.$
In view of the above, we have completed the proof. ∎
Now we are in a position to prove Theorem 1.2.
Proof of Theorem 1.2. Recall that $g$ is a generator of $\mathbb{F}_{q}^{\times}$. Then it is clear that
$\det B_{q}(n)=\det\left[\left(1+\frac{s_{j}}{s_{i}}\right)^{n}\right]_{2\leq i,j\leq n}=\det\left[\left(1+g^{2j-2i}\right)^{n}\right]_{1\leq i,j\leq n-1}.$
By the definition (3.1) of the Teichümuller character $\omega_{q}$ of $\mathbb{F}_{q}$, we see that
(4.2) $\det B_{q}(n)=\det\left[\omega_{q}^{-n}\left(1+g^{2j-2i}\right)\right]_{1\leq i,j\leq n-1}\ {\rm mod}\ \mathfrak{p}=\det W_{n}({\bm{v}})\ {\rm mod}\ \mathfrak{p},$
where
${\bm{v}}=\left(\omega_{q}^{-n}\left(1+g^{0}\right),\omega_{q}^{-n}\left(1+g^{2}\right),\cdots,\omega_{q}^{-n}\left(1+g^{2(n-1)}\right)\right),$
and $\mathfrak{p}=p\mathbb{Z}_{p}[\zeta_{q-1}]$ is a prime ideal. Hence we focus on the determinant of the almost circulant matrix $W_{n}({\bm{v}})$. By Lemma 4.1 and Lemma 3.1 we have
(4.3) $\det B_{q}(n)=\det W_{n}({\bm{v}})\ {\rm mod}\ \mathfrak{p}=\frac{1}{n}\sum_{l=0}^{n-1}\prod_{r\in[0,n-1]\setminus\{l\}}\lambda_{r}\ {\rm mod}\ \mathfrak{p},$
where $\lambda_{r}$ is defined by Eq. (4.1). We next consider $\lambda_{r}\ {\rm mod}\ \mathfrak{p}$. Suppose $r\in[1,n-1]$. Then
$J_{q}\left(\omega_{q}^{-n},\omega_{q}^{-(n+r)}\right)=\frac{G_{q}(\omega_{q}^{-n})G_{q}(\omega_{q}^{-(n+r)})}{G_{q}(\omega_{q}^{-r})}.$
By Lemma 3.5 and Lemma 3.3, we obtain
${\rm ord}_{p}\left(J_{q}\left(\omega_{q}^{-n},\omega_{q}^{-(n+r)}\right)\right)=\frac{1}{p-1}\left(s(n)+s(n+r)-s(r)\right)>0.$
This implies that
$J_{q}\left(\omega_{q}^{-n},\omega_{q}^{-(n+r)}\right)\equiv 0\ ({\rm{mod}}\ \mathfrak{p})$
for any $r\in[1,n-1]$. Hence by Eq. (4.1) we have
(4.4) $\lambda_{r}\equiv\frac{(-1)^{r}}{2}J_{q}\left(\omega_{q}^{-n},\omega_{q}^{-r}\right)\ ({\rm{mod}}\ \mathfrak{p})$
whenever $r\in[1,n-1]$.
Case I: $f\geq 2$.
Suppose $r\in[1,n-1]$. Then
$J_{q}\left(\omega_{q}^{-n},\omega_{q}^{-r}\right)=\frac{G_{q}(\omega_{q}^{-n})G_{q}(\omega_{q}^{-r})}{G_{q}(\omega_{q}^{-(n+r)})}.$
By Lemma 3.5 we obtain
(4.5) ${\rm ord}_{p}\left(J_{q}\left(\omega_{q}^{-n},\omega_{q}^{-r}\right)\right)=\frac{1}{p-1}\left(s(n)+s(r)-s(n+r)\right).$
As $q=p^{f}\geq 9$, we have $n-1\geq(p+3)/2$. By this, for $r\in\{(p+1)/2,(p+3)/2\}\subseteq[1,n-1]$, we have
$s(n)+s(r)>s(n+r).$
This, together with Eq (4.5), implies that
$\left|\left\{1\leq r\leq n-1:J_{q}\left(\omega_{q}^{-n},\omega_{q}^{-r}\right)\equiv 0\ ({\rm{mod}}\ \mathfrak{p})\right\}\right|=\left|\left\{1\leq r\leq n-1:\lambda_{r}\equiv 0\ ({\rm{mod}}\ \mathfrak{p})\right\}\right|\geq 2.$
By this, Eq. (4.4) and Eq. (4.3), we see that $B_{q}(n)$ is singular if $f\geq 2$.
Case II: $f=1$.
Note first that, in this case, $\mathbb{Q}_{p}(\zeta_{q-1})=\mathbb{Q}_{p}$ and $\mathfrak{p}=p\mathbb{Z}_{p}$. By Eq. (4.5), for any $r\in[1,n-1]$, we see that
${\rm ord}_{p}\left(J_{p}\left(\omega_{p}^{-n},\omega_{p}^{-r}\right)\right)=\frac{1}{p-1}\left(s(n)+s(r)-s(n+r)\right)=\frac{1}{p-1}(n+r-(n+r))=0.$
Thus, for any $r\in[1,n-1]$,
$\lambda_{r}\equiv\frac{(-1)^{r}}{2}J_{q}\left(\omega_{p}^{-n},\omega_{p}^{-r}\right)\not\equiv 0\ ({\rm{mod}}\ p\mathbb{Z}_{p}),$
that is, $\lambda_{r}$ is a $p$-adic unit. By this and Eq. (4.3), we obtain
(4.6) $\det B_{q}(n)=-2\lambda_{0}\lambda_{1}\cdots\lambda_{n-1}\left(\frac{1}{\lambda_{0}}+\frac{1}{\lambda_{1}}+\cdots+\frac{1}{\lambda_{n-1}}\right)\ {\rm mod}\ p\mathbb{Z}_{p}.$
We next consider $\lambda_{r}\ {\rm mod}\ p\mathbb{Z}_{p}$. Suppose $r\in[1,n-1]$. Then By Lemma 3.5 and Lemma 3.4 one can verify that
$\lambda_{r}$ $\equiv\frac{(-1)^{r}}{2}J_{p}\left(\omega_{p}^{-n},\omega_{p}^{-r}\right)$
$\equiv\frac{(-1)^{r}}{2}\frac{G_{p}(\omega_{p}^{-n})G_{p}(\omega_{p}^{-r})}{G_{p}(\omega_{p}^{-(n+r)})}$
$\equiv\frac{(-1)^{r+1}}{2}\cdot\frac{\Gamma_{p}\left(\frac{n}{p-1}\right)\Gamma_{p}\left(\frac{r}{p-1}\right)}{\Gamma_{p}\left(\frac{n+r}{p-1}\right)}$
$\equiv\frac{(-1)^{r+1}}{2}\cdot\frac{\Gamma_{p}\left(p-n\right)\Gamma_{p}\left(p-r\right)}{\Gamma_{p}\left(p-n-r\right)}$
$\equiv\frac{(-1)^{r}}{2}\frac{(p-n-1)!\cdot(p-r-1)!}{(p-1-n-r)!}$
$\equiv\frac{(-1)^{r+1}}{2}\frac{(n+r)!}{n!\cdot r!}$
$\equiv\frac{(-1)^{r+1}}{2}\binom{n+r}{r}$
$\equiv-\frac{1}{2}\binom{n}{r}\ ({\rm{mod}}\ p\mathbb{Z}_{p}).$
The last congruence follows from
$(-1)^{r}\cdot\binom{n+r}{r}\equiv\binom{n}{r}\ ({\rm{mod}}\ p\mathbb{Z})$
for any $r\in[1,n-1]$.
By the above results and Sury’s identity [15], we obtain
$-2\lambda_{0}\lambda_{1}\cdots\lambda_{n-1}\left(\frac{1}{\lambda_{0}}+\frac{1}{\lambda_{1}}+\cdots+\frac{1}{\lambda_{n-1}}\right)$
$\equiv$ $(-1)^{n-1}\left(\frac{1}{2}\right)^{n-2}\cdot\prod_{r=0}^{n}\binom{n}{r}\cdot\left(3-2\sum_{r=0}^{n}\binom{n}{r}^{-1}\right)$
$\equiv$ $(-1)^{n-1}\left(\frac{1}{2}\right)^{n-2}\cdot\frac{(n!)^{n+1}}{(0!1!\cdots n!)^{2}}\cdot\left(3-\frac{n+1}{2^{n}}\sum_{k=1}^{n+1}\frac{2^{k}}{k}\right)$
$\equiv$ $(-1)^{n-1}\left(\frac{1}{2}\right)^{n-2}\cdot\frac{(n!)^{n+1}}{(0!1!\cdots n!)^{2}}\cdot\left(3-\frac{1}{2^{n+1}}\sum_{k=1}^{n+1}\frac{2^{k}}{k}\right)$
$\equiv$ $(-1)^{n-1}\left(\frac{1}{2}\right)^{n-2}\cdot\frac{(n!)^{n+1}}{(0!1!\cdots n!)^{2}}\cdot\left(1-\frac{1}{2^{n+1}}\sum_{k=1}^{n}\frac{2^{k}}{k}\right)\ ({\rm{mod}}\ p\mathbb{Z}_{p}).$
By this and using Eq. (2.15) and Eq. (2.9), we finally obtain
$-2\lambda_{0}\lambda_{1}\cdots\lambda_{n-1}\left(\frac{1}{\lambda_{0}}+\frac{1}{\lambda_{1}}+\cdots+\frac{1}{\lambda_{n-1}}\right)$
$\equiv$ $(-1)^{n}\left(\frac{1}{2}\right)^{n-2}\cdot\frac{(n!)^{n+1}}{(0!1!\cdots n!)^{2}}\cdot\frac{2(\frac{2}{p})-2P_{p}-p}{p}\ ({\rm{mod}}\ p\mathbb{Z}_{p}).$
In view of the above, we have completed the proof. ∎
5. Proof of Theorem 1.3
We begin with the following lemma.
Lemma 5.1.
Let $\chi_{q}$ be a generator of $\widehat{\mathbb{F}_{q}^{\times}}$. Then for any nontrivial character $\psi\in\widehat{\mathbb{F}_{q}^{\times}}$, the following results hold.
(5.1) $A_{q}=\prod_{r=0}^{q-2}J_{q}(\psi,\chi_{q}^{r})=\frac{G_{q}(\psi)^{q-1}}{q}.$
(5.2) $S_{q}=\sum_{r=0}^{q-2}\frac{1}{J_{q}(\psi,\chi_{q}^{r})}=\frac{1-q}{q}\left(1+\psi(-1)\right).$
(5.3) $T_{q}=\sum_{r=0}^{q-2}\frac{(-1)^{r}}{J_{q}(\psi,\chi_{q}^{r})}=\frac{1-q}{q}(2-\overline{\psi(2)}).$
Proof.
Since $\chi_{q}$ is a generator of $\widehat{\mathbb{F}_{q}^{\times}}$ and $\psi\neq\varepsilon$, there is a unique integer $j\in[1,q-2]$ such that $\chi_{q}^{-j}=\psi$. One can verify that
$\prod_{r=0}^{q-2}J_{q}(\psi,\chi_{q}^{r})$ $=\prod_{r=0}^{q-2}J_{q}(\chi_{q}^{-j},\chi_{q}^{r})$
$=J_{q}(\chi_{q}^{-j},\chi_{q}^{j})\cdot\prod_{r\in[0,q-2]\setminus\{j\}}J_{q}(\chi_{q}^{-j},\chi_{q}^{r})$
$=(-1)^{j+1}\cdot\prod_{r\in[0,q-2]\setminus\{j\}}\frac{G_{q}(\chi_{q}^{-j})G_{q}(\chi_{q}^{r})}{G_{q}(\chi_{q}^{-j+r})}$
$=(-1)^{j+1}\cdot\frac{G_{q}(\varepsilon)}{G_{q}(\chi_{q}^{-j})G_{q}(\chi_{q}^{j})}\cdot\prod_{r\in[0,q-2]}\frac{G_{q}(\chi_{q}^{-j})G_{q}(\chi_{q}^{r})}{G_{q}(\chi_{q}^{-j+r})}$
$=\frac{G_{q}(\chi_{q}^{-j})^{q-1}}{q}\cdot\prod_{r\in[0,q-2]}\frac{G_{q}(\chi_{q}^{r})}{G_{q}(\chi_{q}^{-j+r})}$
$=\frac{G_{q}(\psi)^{q-1}}{q}.$
This completes the proof of Eq. (5.1).
Now we prove Eq. (5.2). As $|J_{q}(\chi_{q}^{-j},\chi_{q}^{r})|=\sqrt{q}$ whenever $r\in[1,q-2]\setminus\{j\}$, one can verify that
$\sum_{r=0}^{q-2}\frac{1}{J_{q}(\psi,\chi_{q}^{r})}$ $=\sum_{r=0}^{q-2}\frac{1}{J_{q}(\chi_{q}^{-j},\chi_{q}^{r})}$
$=\frac{1}{J_{q}(\chi_{q}^{-j},\varepsilon)}+\frac{1}{J_{q}(\chi_{q}^{-j},\chi_{q}^{j})}+\frac{1}{q}\sum_{r\in[1,q-2]\setminus\{j\}}\frac{q}{J_{q}(\chi_{q}^{-j},\chi_{q}^{r})}$
$=-(1+(-1)^{j})+\frac{1}{q}\sum_{r\in[1,q-2]\setminus\{j\}}\overline{J_{q}(\chi_{q}^{-j},\chi_{q}^{r})}$
$=-(1+\psi(-1))+\frac{1}{q}\sum_{r\in[0,q-2]}\overline{J_{q}(\chi_{q}^{-j},\chi_{q}^{r})}-\frac{1}{q}\overline{J_{q}(\psi,\varepsilon)}-\frac{1}{q}\overline{J_{q}(\psi,\psi^{-1})}$
$=\frac{1-q}{q}\left(1+\psi(-1)\right),$
where the last equality follows from
$\sum_{r\in[0,q-2]}J_{q}(\chi_{q}^{-j},\chi_{q}^{r})=0.$
This completes the proof of Eq. (5.2)
Finally, we prove Eq. (5.3). We first claim that
(5.4) $\sum_{r\in[0,(q-3)/2]}J_{q}(\psi,\chi_{q}^{2r})=\frac{(q-1)\psi(2)}{2}.$
In fact, it is easy to verify that
$\sum_{r\in[0,(q-3)/2]}J_{q}(\psi,\chi_{q}^{2r})$ $=\sum_{r\in[0,(q-3)/2]}\sum_{x\in\mathbb{F}_{q}}\psi(1-x)\chi_{q}^{2r}(x)$
$=\sum_{x\in\mathbb{F}_{q}}\psi(1-x)\sum_{r\in[0,(q-3)/2]}\chi_{q}^{2r}(x)$
$=\frac{q-1}{2}\psi(2),$
where the last equality follows from
$\sum_{r\in[0,(q-3)/2]}\chi_{q}^{2r}(x)=\begin{cases}(q-1)/2&\mbox{if}\ x=\pm 1,\\ 0&\mbox{otherwise.}\end{cases}$
By Eq. (5.4) we obtain
$S_{q}+T_{q}$ $=2\sum_{r\in[0,(q-3)/2]}\frac{1}{J_{q}(\psi,\chi_{q}^{2r})}$
$=\frac{2}{J_{q}(\psi,\varepsilon)}+\frac{1+\psi(-1)}{J_{q}(\psi,\psi^{-1})}+\frac{2}{q}\sum_{\begin{subarray}{c}r\in[1,(q-3)/2]\\ \chi_{q}^{2r}\neq\psi^{-1}\end{subarray}}\frac{q}{J_{q}(\psi,\chi_{q}^{2r})}$
$=-(3+\psi(-1))+\frac{2}{q}\sum_{\begin{subarray}{c}r\in[1,(q-3)/2]\\ \chi_{q}^{2r}\neq\psi^{-1}\end{subarray}}\overline{{J_{q}(\psi,\chi_{q}^{2r})}}$
$=-(3+\psi(-1))+\frac{2}{q}\sum_{r\in[0,(q-3)/2]}\overline{{J_{q}(\psi,\chi_{q}^{2r})}}-\frac{2}{q}\overline{J_{q}(\psi,\varepsilon)}-\frac{1+\psi(-1)}{q}\overline{J_{q}(\psi,\psi^{-1})}$
$=\frac{1-q}{q}\left(3+\psi(-1)-\overline{\psi(2)}\right).$
Combining this with Eq. (5.2), we have
$T_{q}=(S_{q}+T_{q})-S_{q}=\frac{1-q}{q}(2-\overline{\psi(2)}).$
In view of the above, we have completed the proof. ∎
Remark 5.1.
In 1987, Greene [7, Definition 2.4] used Jacobi sums to obtain a finite field analogue of binomial coefficients. In fact, for any $A,B\in\widehat{\mathbb{F}_{q}^{\times}}$, Green defined
$\binom{A}{B}:=\frac{B(-1)}{q}J_{q}(A,\overline{B}).$
By Eq. (5.3) for any nontrivial character $\psi$, we obtain
(5.5) $\sum_{r=0}^{q-2}\binom{\psi}{\chi_{q}^{r}}^{-1}=q\sum_{r=0}^{q-2}\frac{(-1)^{r}}{J_{q}(\psi,\chi_{q}^{-r})}=qT_{q}=(1-q)(2-\overline{\psi(2)}).$
The identity (5.5) can be viewed as a finite field analogue of Sury’s identity
$\sum_{r=0}^{n-1}\binom{n-1}{r}^{-1}=\frac{n}{2^{n}}\sum_{k=1}^{n}\frac{2^{k}}{k}.$
From now on, we fix a generator $g$ of $\mathbb{F}_{q}^{\times}$. For any nontrivial character $\psi\in\widehat{\mathbb{F}_{q}^{\times}}$, define the circulant matrices
(5.6) $M_{q}(\psi):=\left[\psi\left(g^{j-i}-1\right)\right]_{0\leq i,j\leq q-2},$
and
(5.7) $N_{q}(\psi):=\left[\psi\left(g^{j-i}+1\right)\right]_{0\leq i,j\leq q-2}.$
We need the following lemma.
Lemma 5.2.
Let $\chi_{q}$ be a generator of $\widehat{\mathbb{F}_{q}^{\times}}$. Then the numbers
$\alpha_{r}=\psi(-1)J_{q}(\psi,\chi_{q}^{r})\ (r=0,1,\cdots,q-2)$
are exactly all the eigenvalues of $M_{q}(\psi)$. Also, the numbers
$\beta_{r}=(-1)^{r}J_{q}(\psi,\chi_{q}^{r})\ (r=0,1,\cdots,q-2)$
are exactly all the eigenvalues of $N_{q}(\psi)$.
Proof.
For any $r\in[0,q-2]$, it is easy to see that
$\sum_{j=0}^{q-2}\psi(g^{j-i}-1)\chi_{q}^{r}(g^{j})$ $=\sum_{j=0}^{q-2}\psi(g^{j-i}-1)\chi_{q}^{r}(g^{j-i})\chi_{q}^{r}(g^{i})$
$=\sum_{j=0}^{q-2}\psi(g^{j}-1)\chi_{q}^{r}(g^{j})\chi_{q}^{r}(g^{i})$
$=\sum_{x\in\mathbb{F}_{q}}\psi(x-1)\chi_{q}^{r}(x)\chi_{q}^{r}(g^{i})$
$=\psi(-1)J_{q}(\psi,\chi_{q}^{r})\chi_{q}^{r}(g^{i}).$
This implies that
$M_{q}(\psi){\bm{\xi}}_{r}=\alpha_{r}{\bm{\xi}}_{r},$
where
${\bm{\xi}}_{r}=\left(\chi_{q}^{r}(g^{0}),\chi_{q}^{r}(g^{1}),\cdots,\chi_{q}^{r}(g^{q-2})\right)^{T}.$
As these ${\bm{\xi}}_{r}$ are linearly independent over $\mathbb{C}$, the numbers $\alpha_{0},\alpha_{1},\cdots,\alpha_{q-2}$ are all the eigenvalues of $M_{q}(\psi)$. Using essentially the same method, one can easily obtain that all the eigenvalues of $N_{q}(\psi)$ are $\beta_{0},\beta_{1},\cdots,\beta_{q-2}$.
In view of the above, we have completed the proof. ∎
Now we are in a position to prove our last theorem.
Proof of Theorem 1.3. (i) It is clear that
$G(t)=\frac{t^{q-1}-1}{t-1}=1+t+\cdots+t^{q-2}=\prod_{i=2}^{q-1}\left(t-x_{i}\right)$
over $\mathbb{F}_{q}$. This implies
(5.8) $\prod_{i=2}^{q-2}x_{i}=-1.$
Recall that $g$ is a generator of $\mathbb{F}_{q}^{\times}$. By Eq. (5.8) we obtain
(5.9) $\det D_{q}^{-}(\psi)=\prod_{i=2}^{q-1}\psi(x_{i})\cdot\det\left[\psi\left(\frac{x_{j}}{x_{i}}-1\right)\right]_{2\leq i,j\leq q-1}=\psi(-1)\cdot\det\left[\psi(g^{j-i}-1)\right]_{1\leq i,j\leq q-2}.$
Note that $\left[\psi(g^{j-i}-1)\right]_{1\leq i,j\leq q-2}$ is an almost circulant matrix. By Lemma 3.1 and Lemma 5.2, we obtain
$\det\left[\psi(g^{j-i}-1)\right]_{1\leq i,j\leq q-2}$ $=\frac{1}{q-1}\cdot\alpha_{0}\alpha_{1}\cdots\alpha_{q-2}\cdot\left(\frac{1}{\alpha_{0}}+\frac{1}{\alpha_{1}}+\cdots+\frac{1}{\alpha_{q-2}}\right)$
$=\frac{\psi(-1)}{q-1}\cdot\prod_{r=0}^{q-2}J_{q}(\psi,\chi_{q}^{r})\cdot\left(\sum_{r=0}^{q-2}\frac{1}{J_{q}(\psi,\chi_{q}^{r})}\right)$
$=-\frac{(1+\psi(-1))}{q^{2}}G_{q}(\psi)^{q-1}.$
Combining this with Eq. (5.9), we obtain
$\det D_{q}^{-}(\psi)=-\frac{(1+\psi(-1))}{q^{2}}G_{q}(\psi)^{q-1}.$
This completes the proof of (i).
(ii) Similar to Eq. (5.9), we have
(5.10) $\det D_{q}^{+}(\psi)=\psi(-1)\cdot\det\left[\psi(g^{j-i}+1)\right]_{1\leq i,j\leq q-2}.$
Since $\left[\psi(g^{j-i}+1)\right]_{1\leq i,j\leq q-2}$ is an almost circulant matrix, by Lemma 3.1 and Lemma 5.2, one can verify that
$\det\left[\psi(g^{j-i}+1)\right]_{1\leq i,j\leq q-2}$ $=\frac{1}{q-1}\cdot\beta_{0}\beta_{1}\cdots\beta_{q-2}\cdot\left(\frac{1}{\beta_{0}}+\frac{1}{\beta_{1}}+\cdots+\frac{1}{\beta_{q-2}}\right)$
$=\frac{(-1)^{(q-1)/2}}{q-1}\cdot\prod_{r=0}^{q-2}J_{q}(\psi,\chi_{q}^{r})\cdot\left(\sum_{r=0}^{q-2}\frac{(-1)^{r}}{J_{q}(\psi,\chi_{q}^{r})}\right)$
$=\frac{(-1)^{(q+1)/2}}{q^{2}}\left(2-\overline{\psi(2)}\right)G_{q}(\psi)^{q-1}.$
By this and Eq. (5.10), we finally obtain
$\det D_{q}^{+}(\psi)=\frac{(-1)^{(q+1)/2}\cdot\psi(-1)}{q^{2}}\left(2-\overline{\psi(2)}\right)G_{q}(\psi)^{q-1}.$
In view of the above, we have completed the proof. ∎
Acknowledgments The authors would like to thank the referee for helpful comments. We also thank Prof. Hao Pan for his helpful comments and steadfast encouragement.
This research was supported by the Natural Science Foundation of China (Grant Nos. 12101321, 12201291 and 12371004). The first author was supported by the Natural Science Foundation of the Higher Education Institutions of Jiangsu Province (Grant No. 25KJB110010).